# Norm of the Cesàro operator between some spaces of analytic functions

### v1

@@ 1. Introduction
Let $(a)=\{a_{k}\}_{k=0}^{\infty}$ be a sequence of complex numbers. The classical Cesàro operator acting on such sequences is defined as
$\mathcal{C}((a)):=\left(\frac{1}{n+1}\sum_{k=0}^{n}a_{k}\right)_{n=0}^{\infty}.$
As early as the 1920s, Hardy [8] and Landau [13] established the boundedness of this operator on the $\ell^{p}$ spaces.
The Cesàro operator can also be interpreted as an operator acting on functions defined on $\mathbb{D}$, the unit open disc in the complex plane $\mathbb{C}$. Specifically, for any analytic function $f\in H(\mathbb{D})$, if its Taylor expansion is $f(z)=\sum_{k=0}^{\infty}a_{k}z^{k},\ z\in\mathbb{D}$, then the Cesàro operator $\mathcal{C}:H(\mathbb{D})\to H(\mathbb{D})$ is given by
@@
The boundedness and compactness properties of the Cesàro operator have been extensively studied in complex and functional analysis [6, 11, 15, 23]. Early investigations into the boundedness of Cesàro operators on Hardy spaces $H^{p}$ for $1<p<\infty$ relied on Hardy’s work on Fourier series [10] and M. Riesz’s theorem on conjugate functions [7, Theorem 4.1]. Using the theory of composition operator semigroups, Siskakis [21] gave an alternative proof of the above result and further studied the case $p=1$ in [20]. A different proof for the case $p=1$ was provided by Giang [14]. Later, Miao [15] showed that the Cesàro operator is bounded on $H^{p}$ for all $0<p<1$.
@@ 2. Notation Preliminaries
$\frac{3}{2}$ $\leq\|\mathcal{C}\|_{H^{\infty}\rightarrow\mathcal{B}^{\alpha}}\leq 1+3\cdot 2^{\alpha-1}~~~\mbox{if}~~~\alpha>1.$
@@
$\|\mathcal{C}f\|_{\mathcal{B}^{\alpha}}\leq 1+\sup_{0<r<1}h(r)(1-r^{2})^{\alpha-1}\leq 1+3\cdot 2^{\alpha-1}.$

### v2

@@ 1. Introduction
Let $(a)=\{a_{k}\}_{k=0}^{\infty}$ denote a sequence of complex numbers. The classical Cesàro operator, which acts on sequences of this form, is defined by the relation
$\mathcal{C}((a)):=\left(\frac{1}{n+1}\sum{k=0}^{n}a_{k}\right)_{n=0}^{\infty}.$
The boundedness of this operator on the $\ell^{p}$ spaces was already established in the 1920s, through the work of Hardy [8] and Landau [13].
An alternative interpretation of the Cesàro operator is as an operator on functions defined on the unit open disc $\mathbb{D}\subset\mathbb{C}$. More precisely, for an analytic function $f\in H(\mathbb{D})$ with Taylor expansion $f(z)=\sum_{k=0}^{\infty}a_{k}z^{k}$ ($z\in\mathbb{D}$), the Cesàro operator $\mathcal{C}:H(\mathbb{D})\to H(\mathbb{D})$ is defined by
@@
The boundedness and compactness of the Cesàro operator have been the subject of extensive research in complex and functional analysis, as documented in [6, 11, 15, 23]. Initial studies of its boundedness on Hardy spaces $H^{p}$ for $1<p<\infty$ drew on Hardy’s results concerning Fourier series [10] and M. Riesz’s theorem on conjugate functions [7, Theorem 4.1]. By employing the theory of composition operator semigroups, Siskakis [21] provided an alternative demonstration of this boundedness and extended the investigation to the case $p=1$ in [20]. A separate proof for the $p=1$ case was later contributed by Giang [14]. Subsequently, Miao [15] established that the Cesàro operator remains bounded on $H^{p}$ for all exponents $0<p<1$.
@@ 2. Notation Preliminaries
$\frac{3}{2}$ $\leq\|\mathcal{C}\|_{H^{\infty}\rightarrow\mathcal{B}^{\alpha}}\leq 4~~~\mbox{if}~~~\alpha>1.$
@@
$\|\mathcal{C}f\|_{\mathcal{B}^{\alpha}}\leq 1+\sup_{0<r<1}h(r)(1-r^{2})^{\alpha-1}\leq 1+3\sup_{0<r<1}(1-r^{2})^{\alpha-1}=4.$